# spacing=nonfrench

### v2

@@ Abstract
This paper introduces SysSpec, a framework for developing generative file systems. Its key insight is to replace ambiguous natural language with principles adapted from formal methods. Instead of imprecise prompts, SysSpec employs a multi-part specification that accurately describes a file system’s functionality, modularity, and concurrency. The specification acts as an unambiguous blueprint, guiding LLMs to generate expected code flexibly. To manage evolution, we develop a DAG-structured patch that operates on the specification itself, enabling new features to be added without violating existing invariants. Moreover, the SysSpec toolchain features a set of LLM-based agents with mechanisms to mitigate hallucination during construction and evolution. We demonstrate our approach by generating SpecFS, a concurrent file system. SpecFS passes hundreds of regression tests, matching a manually-coded baseline. We further confirm its evolvability by seamlessly integrating 10 real-world features from Ext4. Our work shows that a specification-guided approach makes generating and evolving complex systems not only feasible but also highly effective.
footnotetext: ✉ Co-corresponding authors.
@@ 1 Introduction
Nevertheless, although recent advancements in LLMs have demonstrated their profound capabilities in automated code generation [dong2025qimeng, qi2025intention, cursor, copilot, zhang2024codeagentenhancingcodegeneration, huang2024agentcodermultiagentbasedcodegeneration, wang2025coderagbenchretrievalaugmentcode, zhang2024autocoderover], generating a complete and useful file system from natural-language prompts is profoundly challenging, if not impossible. The intricate semantics of file systems, from ensuring concurrency correctness to carefully managing complex file structures, are difficult to express unambiguously in prompts. Consequently, to our knowledge, no prior work has successfully generated a complete, functional, and feature-rich file system using an LLM-based approach.
@@
To our knowledge, SysSpec is the first framework that enables both the generation and principled evolution of end-to-end file systems like SpecFS, shifting the developer’s focus from writing brittle low-level code to crafting robust, high-level designs. SysSpec and SpecFS will be open-sourced.
@@ 2.1 The Anatomy of File System Change
Methodology. We choose Ext4 as our target because it is a mature and (still) widely-deployed file system with 20 years of evolution in real-world environments. Ext4 was introduced in Linux 2.6.19 and has been continuously developed, incorporating thousands of patches related to performance, new features, and maintenance. Our analysis is based on all 3,157 Ext4-related commits merged into the mainline Linux from version 2.6.19 to 6.15. To understand the evolution, we categorize each commit using a classification scheme for FS patches adapted from prior work [evolutionLu2014]: (1) Bug: fixing an existing bug, (2) Performance: improving efficiency through new designs or optimizations, (3) Reliability: enhancing the file system’s robustness, (4) Feature: implementing new functionalities, and (5) Maintenance: refactoring code or improving documentation without changing semantic behavior.
@@
Implication-4: Evolution is taken in small steps. We measured the lines of code (LOC) for each patch and present the cumulative distribution function (CDF) in Fig. 3. Our analysis shows that evolution proceeds through small, frequent changes. Specifically, about 80% of all bug fixes involve fewer than 20 LOC. While feature-related patches are generally larger, about 60% of them still require fewer than 100 LOC. Moreover, the vast majority of commits also modify only a single file, as shown in Fig. 2-b. These findings suggest that file system evolution is composed of these manageable, small-scale changes.
@@ 2.2 Case Study: Evolution of Fast Commit
-- if (unlikely(error))
++ if (unlikely(error)) {
++ ext4_fc_stop_update(inode);
@@ 3 Design Overview
The SysSpec toolchain. SysSpec provides an LLM-based toolchain that translates the high-level specification into an executable implementation. As illustrated in Fig. 5-b, this toolchain includes agents: The SpecCompiler agent translates the specification (SpecFS) into a low-level C-language implementation (ImpFS). To manage the current limitations of LLMs, it uses techniques like two-phase prompting to handle complex logic. The SpecValidator agent ensures the correctness of the generated code, employing a retry-with-feedback loop to automatically identify and correct errors. With the designs of SysSpec, our evaluation shows achieves significant correctness improvements for complex operations compared to naive prompting (up to 34.4%). We also provide SpecAssistant to ease the development of specification.
SpecFS. With SysSpec, we design and implement SpecFS, a complete file system written entirely in the SysSpec specification language. Our toolchain can automatically generate a functional C-language implementation of SpecFS without human intervention. Through a series of case studies, we demonstrate that SpecFS can seamlessly evolve via spec patches to support sophisticated features found in Ext4, e.g., extents, delayed allocation, and file encryption.
@@ 4.1 Functionality Specification
The functionality specification defines the behavior of a single module by describing its state transitions. A module, which we define in §4.2, is a collection of related state variables and functions. The specification is built upon three components, inspired by formal methods.
First, pre- and post-conditions, following Hoare logic, define the contractual obligations for each function by specifying the required state before execution and the guaranteed state upon completion. Second, invariants are properties that must hold true across all state transitions, ensuring the module’s integrity. Finally, a system algorithm outlines the high-level logic for how a function should achieve its state transition, guiding the LLM’s implementation strategy. Rather than always providing a complete system algorithm, our experience shows that a high-level intent is often sufficient. The intent can be regarded as a lightweight system algorithm that, expressed in natural language, guides the LLM in generating the desired implementation.
@@
Post-condition:
Case 1 Successful traversal and insertion
@@
Invariant: root_inum always exists
Figure 6: Simplified functionality specification for atomfs_ins(path,name)
Hoare logic for file system synthesis. To specify function behavior, we adapt the classic Hoare logic formalism of {P}C{Q} using pre- and post-conditions, but we sidestep the high complexity of full formal verification. Each function is annotated with pre-conditions that define required system states and post-conditions that guarantee specific state transitions and return values. Our approach diverges from traditional formal methods in two key aspects. First, our specifications are expressed in structured natural language augmented with type annotations, rather than pure mathematical logic, making them more accessible. Second, an LLM enforces adherence to this logic during code generation, replacing the role of a formal theorem prover.
Fig. 6 shows atomfs_ins, an internal function of AtomFS [zou2019atomfs] that implements mknod/mkdir. The pre-condition describes the validness of the parameters. The post-condition describes the state after the operation.
@@ 4.2 Modularity Specification
SpecFS addresses the composition challenge through a methodology that combines interface contracts with LLM context management.
Context-bounded modular synthesis. One of the core innovations in SysSpec’s modularity specification lies in aligning module design with two intrinsic properties of LLM reasoning. First, strict size constraints ensure each module ($\leq$500 LoC) fits entirely within the model’s context window, enabling holistic analysis during generation. Second, explicit interface contracts govern inter-module dependencies through Rely-Guarantee conditions, a formal mechanism adapted from concurrent program verification [zou2019atomfs]. This methodology embodies three critical principles: (1) module implementations must respect their declared dependencies (Rely), (2) provide guarantees about their behavior (Guarantee), and (3) compose through logical implication of these contracts.
[RELY]
Predefined Structures/Functions:
struct inode { ... };
@@
void insert(struct inode*, struct inode*, char*) // ...
int check_ins(struct inode*, char*) // ...
@@
[Rely]
[Locking Specifications of locate]
Pre-condition: cur is locked.
Post-condition: suppose return target.
@@
[Locking Specifications of check_ins]
Pre-condition: cur is locked.
Post-condition:
@@
Pre-condition: no lock is owned.
Post-condition: no lock is owned.
@@ 4.3 Concurrency Specification
int atomfs_ins(char* path[], char* name, ...) {
@@
struct inode* new_inode = malloc_inode(...);
@@ 4.4 DAG-Structured Specification Patch
Intermediate nodes: building on guarantees. An intermediate node represents a step in the evolution that builds upon previously introduced changes. Its specification relies on the new $\mathsf{guarantee}$s provided by its child nodes to implement more complex logic. In turn, this node provides its own, higher-level guarantees, forming a clear dependency chain that progressively constructs the new feature.
The root node: the integration point. The root node is the culmination of the evolution, acting as the final integration point. Its specification provides a new $\mathsf{guarantee}$ that is semantically equivalent to an existing guarantee in the original system. This equivalence is critical, as it allows the entire chain of new functionality, built up through the DAG, to safely and transparently replace an old implementation. This substitution serves as the “commit point”, where the evolution is atomically applied to the base system.
The evolution process. The unidirectional dependencies between nodes naturally form a DAG, which dictates the evolution workflow. The process begins with the SysSpec toolchain generating code for the leaf nodes. It then traverses the graph upwards, synthesizing the implementation for each parent node by leveraging the freshly generated guarantees from its children. This continues until the toolchain reaches the root node, at which point the new feature is fully implemented and integrated into the file system.
@@ 4.5 The SysSpec Toolchain
We have designed and implemented three LLM-based agents, SpecCompiler, SpecValidator, and SpecAssistant, that form the core toolchain for the SysSpec framework.
@@
The SpecValidator agent. The SpecValidator agent performs the final, holistic verification of the complete ImpFS. It adopts a two-pronged strategy that combines specification-based review with traditional testing. First, it re-uses the SpecEval logic from the SpecCompiler to check each fully-generated module against the combined functionality and concurrency specifications. Second, it integrates with standard software engineering workflows by running a suite of unit and regression tests against the final C-language file system. This process emulates a modern CI/CD pipeline: the SpecEval component acts as an automated code reviewer verifying adherence to the design, while the test suite ensures that no existing functionality has regressed.
The SpecAssistant agent. The SpecAssistant agent is an interactive assistant that streamlines specification development through a human-in-the-loop process. A developer provides a draft specification, which the SpecAssistant first validates and reformats to meet SysSpec’s syntax. The agent then enters an automated refinement loop. It repeatedly invokes the SpecCompiler; if the SpecCompiler’s SpecEval phase identifies a flaw, the SpecAssistant uses a new SpecFine step to automatically polish the specification based on the feedback before retrying. This loop concludes in one of two ways. On success, the SpecAssistant provides the developer with the refined specification and the generated C implementation for validation. On failure, it returns the last attempted specification annotated with detailed diagnostics, serving as a debug log that guides the developer in resolving the issue.
@@ 5.1 Prototype Implementation
SpecFS is organized into 45 distinct modules distributed across several logical layers, including file operations, inode management, path traversal, and the POSIX interface. SpecFS supports a wide range of standard POSIX calls, such as $\mathsf{open}$, $\mathsf{read}$, and $\mathsf{rename}$. To validate its functional correctness, we use the xfstests [xfstest] suite within our SpecValidator agent. SpecFS demonstrates a level of correctness comparable to that of the original AtomFS, failing only 64 out of 754 test cases, all attributable to unimplemented functionality.
System complexity. While SpecFS serves as our prototype, it is a non-trivial file system by any measure. The generated C implementation comprises approximately 4,300 lines of code. To put this figure into context, we compared it against the 82 file systems in the Linux 6.1.10 kernel. SpecFS ranks 42nd by line count, surpassing established systems like squashfs and nearly matching the size of 9pfs.
@@
Figure 10: DAG-Structured patch for implementing “Extent”. For simplicity, a node in the figure may encompass multiple modules of specifications. “Common” denotes modifications made to some existing data structures, serving as the “rely” within other nodes’ corresponding specifications. “Root Identifier” indicates that this node is the root (along with its associated logic.)
@@ 5.2 Case Study of Evolutions
Example: Extent. This feature transforms the file structure from individual blocks into extents. Each extent records a segment of contiguous blocks, facilitating sequential reads and writes. The modification process (Fig. 10) begins by altering the low-level file operations in lowlevel_file (and corresponding initializations) to incorporate new extent-based file I/O operations. Subsequently, the inode_management module is modified to invoke these new extent operations, ultimately serving as the root node, making the complete set of new operations visible to the entire system.

### v3

@@ Abstract
This paper introduces SysSpec, a framework for developing generative file systems. Its key insight is to replace ambiguous natural language with principles adapted from formal methods. Instead of imprecise prompts, SysSpec employs a multi-part specification that accurately describes a file system’s functionality, modularity, and concurrency. The specification acts as an unambiguous blueprint, guiding LLMs to generate expected code flexibly. To manage evolution, we develop a DAG-structured patch that operates on the specification itself, enabling new features to be added without violating existing invariants. Moreover, the SysSpec toolchain features a set of LLM-based agents with mechanisms to mitigate hallucination during construction and evolution. We demonstrate our approach by generating SpecFS, a concurrent file system. SpecFS demonstrates equivalent level of correctness to that of a manually-coded baseline across hundreds of regression tests. We further confirm its evolvability by seamlessly integrating 10 real-world features from Ext4. Our work shows that a specification-guided approach makes generating and evolving complex systems not only feasible but also highly effective.
@@ 1 Introduction
Nevertheless, although recent advancements in LLMs have demonstrated their profound capabilities in automated code generation [dong2025qimeng, cursor, copilot, qi2025intention, zhang2024codeagentenhancingcodegeneration, huang2024agentcodermultiagentbasedcodegeneration, wang2025coderagbenchretrievalaugmentcode, zhang2024autocoderover], generating a complete and useful file system from natural-language prompts is profoundly challenging, if not impossible. The intricate semantics of file systems, from ensuring concurrency correctness to carefully managing complex file structures, are difficult to express unambiguously in prompts. Consequently, to our knowledge, no prior LLM-based approach has successfully generated a complete, functional, and feature-rich file system.
@@
To our knowledge, SysSpec is the first framework that enables both the generation and principled evolution of end-to-end file systems like SpecFS, shifting the developer’s focus from writing brittle low-level code to crafting robust, high-level designs. SysSpec and SpecFS is available at: https://llmnativeos.github.io/specfs/.
@@ 2.1 The Anatomy of File System Change
Methodology. We choose Ext4 as our target because it is a mature and (still) widely-deployed file system with 20 years of evolution in real-world environments. Ext4 was introduced in Linux 2.6.19 and has been continuously developed, incorporating thousands of patches related to performance, new features, and maintenance. Our analysis is based on all 3,157 Ext4-related commits (i.e., individual patches rather than patch sets) merged into the mainline Linux from version 2.6.19 to 6.15. To understand the evolution, we categorize each commit using a classification scheme for FS patches adapted from prior work [evolutionLu2014]: (1) Bug: fixing an existing bug, (2) Performance: improving efficiency through new designs or optimizations, (3) Reliability: enhancing the file system’s robustness, (4) Feature: implementing new functionalities, and (5) Maintenance: refactoring code or improving documentation without changing semantic behavior.
@@
Implication-4: Evolution is taken in small steps. We measured the LOC for each patch and present the cumulative distribution function (CDF) in Fig. 3. Our analysis shows that evolution proceeds through small, frequent changes. Specifically, about 80% of all bug fixes involve fewer than 20 LOC. While feature-related patches are generally larger, about 60% of them still require fewer than 100 LOC. Moreover, the vast majority of commits also modify only a single file, as shown in Fig. 2-b. These findings suggest that file system evolution is composed of these manageable, small-scale changes.
@@ 2.2 Case Study: Evolution of Fast Commit
[…]
@@ 3 Design Overview
The SysSpec toolchain. SysSpec provides an LLM-based toolchain that translates the high-level specification into an executable implementation. As illustrated in Fig. 5-b, this toolchain includes agents: The SpecCompiler agent translates the specification (SpecFS) into a low-level C-language implementation (ImpFS). To manage the current limitations of LLMs, it uses techniques like two-phase prompting to handle complex logic. The SpecValidator agent ensures the correctness of the generated code, employing a retry-with-feedback loop to automatically identify and correct errors. With the designs of SysSpec, our evaluation shows significant correctness improvements for complex operations compared to naive prompting (up to 34.4%). We also provide SpecAssistant to ease the development of specification.
Case study: SpecFS. With SysSpec, we design and implement SpecFS, a complete file system written entirely in the SysSpec specification language. Our toolchain can automatically generate a functional C-language implementation of SpecFS without human intervention. Through a series of case studies, we demonstrate that SpecFS can seamlessly evolve via spec patches to support sophisticated features found in Ext4, e.g., extents or delayed allocation.
@@ 4.1 Functionality Specification
The functionality specification defines the behavior of a module by describing its state transitions. A module is a collection of related state variables and functions. The specification is built upon three components, inspired by formal methods.
First, pre- and post-conditions, following Hoare Logic, define the contractual obligations for each function by specifying the required state before execution and the guaranteed state upon completion. Second, invariants are properties that must hold true across all state transitions, ensuring the module’s integrity. Such Hoare-logic-based specification approach inherently avoids the need for explicitly constructing the entire state space, thereby directly circumventing the problem of state space explosion. Finally, a system algorithm outlines the high-level logic for how a function should achieve its state transition, guiding the LLM’s implementation strategy. Rather than always providing a complete system algorithm, our experience shows that a high-level intent is often sufficient. The intent can be regarded as a lightweight system algorithm that, expressed in natural language, guides the LLM in generating the desired implementation.
@@
[…]
@@
[…]
Figure 6: Simplified functionality specificaiton for atomfs_ins
Hoare logic for file system synthesis. To specify function behavior, we adapt the classic Hoare logic formalism of {P}C{Q} using pre- and post-conditions, but we sidestep the high complexity of full formal verification. Each function is annotated with pre-conditions that define required system states and post-conditions that guarantee specific state transitions and return values. Fig. 6 shows atomfs_ins, an internal function of AtomFS [zou2019atomfs] that implements mknod/mkdir. The pre-condition describes the validness of the parameters. The post-condition describes the state after the operation.
Our approach diverges from traditional formal methods in two key aspects. First, an LLM enforces adherence to this logic during code generation, replacing the role of a formal theorem prover. Second, to balance the comprehensibility and unambiguity of the specification, our specifications are expressed in structured natural language augmented with type annotations rather than pure mathematical logic, thereby making them more accessible. SysSpec augments the semantic precision of natural language with a structured organization for the specifications (e.g., sections for functionality, modularity, and concurrency), and uses mathematically disciplined natural-language expressions to ensure that the intended semantics remain precise and unambiguous while keeping the specification accessible. For example, the specification states that the file size equals $max(\text{old\_size},\text{offset}+\text{len})$, rather than that the write updates the file size if necessary. These designs significantly limit the potential for misinterpretation.
@@ 4.2 Modularity Specification
SysSpec addresses the composition challenge, avoiding interface-level dependency errors through a methodology that combines interface contracts with LLM context management.
Context-bounded modular synthesis. One of the core innovations in SysSpec’s modularity specification lies in aligning module design with two intrinsic properties of LLM reasoning. First, strict size constraints ensure each module fits entirely within the model’s context window, enabling holistic analysis during generation. The specific constraints on module sizes evolve in tandem with improvements in LLM capabilities and context window capacities. Take our case study (§5) as an example, we limited module sizes to $\leq$500 LoC, which keeps the token consumption for inference generally within approximately 30K tokens. Second, explicit interface contracts govern inter-module dependencies through Rely-Guarantee conditions, a formal mechanism adapted from concurrent program verification [zou2019atomfs]. This methodology embodies three critical principles: (1) module implementations must respect their declared dependencies (Rely), (2) provide guarantees about their behavior (Guarantee), and (3) compose through logical implication of these contracts.
[…]
struct inode { };
@@
void insert(struct inode*, struct inode*, char*) //
int check_ins(struct inode*, char*) //
@@
Incorporation with external code. SysSpec also supports incorporating external code (e.g., libraries) via the Rely-Guarantee framework. External code can be integrated by first exposing their Guarantees. Developers then specify dependencies on these exposed guarantees within the Rely clause of their specifications. During code generation, the external code is treated as a satisfied dependency, enabling the generated module to correctly invoke external functions without reimplementing them.
[…]
@@
[…]
@@
[…]
@@ 4.3 Concurrency Specification
int atomfs_ins(char* path[], char* name, …) {
@@
struct inode* new_inode = malloc_inode(…);
@@ 4.4 DAG-Structured Specification Patch
Intermediate nodes: building on guarantees. An intermediate node represents a step in the evolution that builds upon previously introduced changes. Its specification relies on the new $\mathsf{guarantee}$s provided by its child nodes to implement more complex logic. In turn, this node provides its own, higher-level guarantees, forming a clear dependency chain that progressively constructs the new feature. In principle, any modification to a guarantee necessitates corresponding specification adjustments for all modules that depend on it, i.e., these modules are included in the patch as intermediate nodes to preclude potential compatibility issues.
The root node: the integration point. The root node is the culmination of the evolution, acting as the final integration point. Unlike the “tree” structure, a “DAG” structured patch may have multiple root nodes. Root nodes are characterized by the property that their specification provides semantically unchanged $\mathsf{guarantees}$. This equivalence is critical, as it allows the entire chain of new functionality, built up through the DAG, to safely and transparently replace an old implementation. This substitution serves as the “commit point”, where the evolution is atomically applied to the base system.
Evolution involving existing modules. Nodes within a DAG-structured specification patch includes both new modules and modifications to existing ones. A modified existing module is treated as a “new module,” which can largely reuse the existing specification. If a shared component (e.g., inode) is modified, all dependent modules must be regenerated to “rely” on the updated version, while other parts of their specifications may not require any modifications. These modified modules then replace the original ones after the patch is merged.
The evolution process. The unidirectional dependencies between nodes naturally form a DAG, which dictates the evolution workflow. The process begins with the SysSpec toolchain (§4.5) generating code for the leaf nodes. It then traverses the graph upwards, synthesizing the implementation for each parent node by leveraging the freshly generated guarantees from its children. This continues until the toolchain reaches the root node, at which point the new feature is fully implemented and integrated into the file system.
@@ 4.5 The SysSpec Toolchain
We have designed and implemented three LLM-based agents, SpecCompiler, SpecValidator, and SpecAssistant, that form the core toolchain for the SysSpec to serve the entire lifecycle of generative file system development, The toolchain further effectively mitigates hallucinations in LLMs by addressing two key aspects: (1) it helps that LLMs produce outputs aligned with those specifications through the coordination of SpecCompiler and SpecValidator, and (2) it helps developers in producing correct specifications by leveraging a strictly structured specification format and the SpecAssistant.
@@
The SpecValidator agent. The SpecValidator agent performs the final, holistic verification of the complete ImpFS. It combines specification-based review with traditional testing. First, it re-uses SpecEval logic from SpecCompiler to check each fully-generated module against the combined functionality and concurrency specifications. Second, it integrates with standard software engineering workflows by running a suite of unit and regression tests against the final C-language file system. This process emulates a modern CI/CD pipeline: the SpecEval component acts as an automated code reviewer verifying adherence to the design, while the test suite ensures that no existing functionality has regressed.
The SpecAssistant agent. SpecAssistant streamlines specification development through a human-in-the-loop process. A developer provides a draft specification, which the SpecAssistant first validates and reformats to meet SysSpec’s syntax. The agent then enters an automated refinement loop. It repeatedly invokes the SpecCompiler; if the SpecCompiler’s SpecEval phase identifies a flaw, the SpecAssistant uses a new SpecFine step to automatically polish the specification based on the feedback before retrying. This loop concludes in two ways. On success, the SpecAssistant provides the developer with the refined specification and the C implementation for validation. On failure, it returns the last attempted specification annotated with detailed diagnostics, serving as a debug log that guides the developer in resolving the issue.
@@ 5.1 Prototype Implementation
SpecFS is organized into 45 distinct modules distributed across several logical layers, including file operations, inode management, path traversal, and the POSIX interface. SpecFS supports a wide range of standard POSIX calls, such as $\mathsf{open}$, $\mathsf{read}$, and $\mathsf{rename}$. To validate its functional correctness, we use the xfstests [xfstest] suite within our SpecValidator. SpecFS demonstrates a level of correctness equivalent to that of the original AtomFS, failing only 64 out of 754 test cases, all attributable to unimplemented functionality.
System complexity and comprehensibility. While SpecFS serves as our prototype, it is a non-trivial file system by any measure. The generated C implementation comprises approximately 4,300 lines of code. To put this figure into context, we compared it against the 82 file systems in the Linux 6.1.10 kernel. SpecFS ranks 42nd by line count, surpassing established systems like squashfs and nearly matching the size of 9pfs. We choose C as the programming language as it remains the de facto standard and is most widely used for low-level file system developments. While high-level languages like Rust may be applicable in the future, they present trade-offs between the benefits from their inherent features (e.g., memory safety) and the increased specification complexity (e.g., due to Rust’s ownership model). Moreover, we observe that the generated code of SpecFS is highly comprehensible. LLMs naturally adhere to standard coding conventions and generate explaining comments, mirroring the style of human engineers.
@@
Experience for developing SpecFS. We gain several experiences during the development of SpecFS. First, while strict SysSpec adherence effectively ensures accuracy of code generation (§6.1), simplified specifications (e.g., simplify Rely/Guarantee) can occasionally still result in logically correct code. If the objective is to maximize development efficiency, one might consider making a trade-off between generation accuracy and the use of simplified specifications. Second, to debug SpecFS, we initially locate bugs in the generated C code, consistent with debugging methods for human-written code. Once a bug is localized, the debugging process additionally involves validating the high-level design against the specification. In our experience, most issues in SpecFS are diagnosed at the specification level.
@@ 5.2 Case Study of Evolutions
Example: Extent. This feature transforms the file structure from individual blocks into extents. Each extent records a segment of contiguous blocks, facilitating sequential reads and writes. The modification process (Fig. 10) begins by defining new data structure for extent and inode. Then, it updates the low-level file operations in lowlevel_file (and corresponding initializations) to incorporate new extent-based file I/O operations. Subsequently, the inode_management module is modified to invoke these new extent operations. Since the new inode_management provides the same guarantee as the original inode_management, the new inode_management serves as the root node of the entire patch and directly replaces the original one, making the complete set of new operations visible to the entire system.
Figure 10: DAG-Structured patch for implementing “Extent”. For simplicity, a node in the figure may encompass multiple modules of specifications. “Root Identifier” indicates that this node is the root (along with its associated logic).